The rank of elliptic curves over real quadratic number fields

of class number 1

Dissertation zur Erlangung des Grades

des Doktors der Naturwissenschaften der Mathematisch–Naturwissenschaftlichen Fakultät

der Universität des Saarlandes

von Pascale Serf

Saarbrücken

Tag des Kolloquiums: 13. Dezember 1995
Dekan:  Prof. Dr. H. Becker
Berichterstatter:  Prof. Dr. H.G. Zimmer

Dr. J.E. Cremona, Exeter

Contents

Preface v

Deutsche Zusammenfassung ix

List of notation xiii

1 Algebraic number theory preliminaries 1 1.1 A special case of Hensel’s lemma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Quadratic number fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

1.2.1 Field discriminant and integral basis . . . . . . . . . . . . . . . . . . . . . . . . 4 1.2.2 Unit group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 1.2.3 Decomposition law of rational primes . . . . . . . . . . . . . . . . . . . . . . . 5 1.2.4 A set of representatives of the residue class field . . . . . . . . . . . . . . . . . 6 1.2.5 The completion K_{P} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 1.2.6 P-adic values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 1.2.7 Squares in K_{P} (“Local squares”) . . . . . . . . . . . . . . . . . . . . . . . . . . 14 1.2.8 “Global squares” and square roots . . . . . . . . . . . . . . . . . . . . . . . . . 16 1.2.9 How to find generators of prime ideals (Pell’s equation) . . . . . . . . . . . . . 18

1.3 Real quadratic number fields of class number 1 . . . . . . . . . . . . . . . . . . . . . . 21 1.3.1 Conditions on D for h(ℚ(\sqrt{D})) to be odd . . . . . . . . . . . . . . . . . . . . . 21 1.3.2 Distribution of class number 1 fields . . . . . . . . . . . . . . . . . . . . . . . . 22 1.3.3 Where do we use class number 1? . . . . . . . . . . . . . . . . . . . . . . . . . . 22

2 2-descent via 2-isogeny 25 2.1 The algorithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 2.2 Local solubility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

2.2.1 Consequences of Hensel’s lemma . . . . . . . . . . . . . . . . . . . . . . . . . . 31 2.2.2 Local solubility modulo a finite prime . . . . . . . . . . . . . . . . . . . . . . . 35 2.2.3 Local solubility modulo the infinite prime(s) . . . . . . . . . . . . . . . . . . . . 40

2.3 Global solubility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 2.4 Points on the elliptic curve from points on the quartics . . . . . . . . . . . . . . . . . . 42

2.4.1 Torsion points . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

2.4.2 Linear independence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 2.5 Tricks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

2.5.1 Sieving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 2.5.2 Using group structures to restrict the number of d_{1} to be considered . . . . . . 48

3 The Tate-Shafarevich group 55 3.1 Some general results . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 3.2 Selmer and Tate-Shafarevich group for the ϕ-descent . . . . . . . . . . . . . . . . . . . 56 3.3 S^{(ϕ)}(E/K) and � (E/K)[ϕ] for elliptic curves with non-trivial 2-torsion . . . . . . . . 58 3.4 � (E/K)[2], the 2-torsion of � (E/K) . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 3.5 Examples of large � (E/ℚ(\sqrt{D}))[2] . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

A Examples of ranks and the 2-torsion of the Tate-Shafarevich group of elliptic curves with a point of order 2 69 A.1 Four curves over ℚ(\sqrt{5}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 A.2 Higher ranks over ℚ(\sqrt{5}) and ℚ(\sqrt{6}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

4 General 2-descent 93 4.1 The principle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 4.2 Finding the quartics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96

4.2.1 Reduced quadratic forms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 4.2.2 Bounding Im(σ_{1}(G)(z_{1})) and Im(σ_{2}(G)(z_{2})) simultaneously . . . . . . . . . . . 99 4.2.3 Bounding the product Im(σ_{1}(G)(z_{1})) · Im(σ_{2}(G)(z_{2})) . . . . . . . . . . . . . . . 102 4.2.4 Reduced quartics ax^{4} + bx^{3} + cx^{2} + dx + e over ℝ and the resulting bounds on

a and c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 4.2.5 Which types of quartics have to be considered? . . . . . . . . . . . . . . . . . . 111 4.2.6 All possible coefficients a of reduced quartics ax^{4}+ bx^{3}+ cx^{2}+ dx + e over 𝔒(K) 112 4.2.7 All possible coefficients b of reduced quartics ax^{4} + bx^{3} + cx^{2} + dx + e over 𝔒(K)

for a fixed a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 4.2.8 All possible coefficients c of reduced quartics ax^{4} + bx^{3} + cx^{2} + dx + e over 𝔒(K)

for a fixed pair (a, b) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 4.2.9 How to compute the coefficients d and e of ax^{4} + bx^{3} + cx^{2} + dx + e for a fixed

triple (a, b, c) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 4.2.10 Summary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 4.2.11 A small problem arising during the computation of d and e . . . . . . . . . . . 123

4.3 Testing triviality . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 4.4 Testing equivalence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 4.5 Local and global solubility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 4.6 Points on the elliptic curve from points on the quartics . . . . . . . . . . . . . . . . . . 136 4.7 Tricks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140

4.7.1 Divisibility conditions in the a, b, c–loop . . . . . . . . . . . . . . . . . . . . . . 140 4.7.2 Sieving . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 5 The relevant (I, J) pairs 145

5.1 … over ℚ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 5.1.1 Lemma 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 5.1.2 Lemma 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 5.1.3 Lemma 5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 5.1.4 One or two (I, J) pairs? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 5.1.5 Divisibility conditions on quartics satisfying 2^{4} | I and 2^{6} | J . . . . . . . . . . 165 5.1.6 Further divisibility conditions on quartics satisfying 2^{6} | I and 2^{7} | J . . . . . . 168 5.1.7 Why 2^{6} | I, 2^{9} | J, 2^{10} | 8I + J in Lemma 5? . . . . . . . . . . . . . . . . . . . 169

5.2 … over number fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 5.2.1 Lemma 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 5.2.2 Lemma 4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 5.2.3 Lemma 5^{∗} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 5.2.4 Divisibility conditions on quartics satisfying π^{4} | I and π^{6} | J . . . . . . . . . . 209 5.2.5 Divisibility conditions on quartics satisfying π^{2ε+2} | I and π^{2ε+4} | J . . . . . . . 212 5.2.6 Further divisibility conditions on quartics satisfying π^{2ε+4} | I and π^{2ε+5} | J . . 214 5.2.7 The divisibility conditions in practice . . . . . . . . . . . . . . . . . . . . . . . 216 5.2.8 Lemma 5 over quadratic number fields . . . . . . . . . . . . . . . . . . . . . . . 218 5.2.9 One, two, three, or four (I, J) pairs over quadratic number fields? . . . . . . . 224

B Examples of ranks of elliptic curves with trivial 2-torsion 229 B.1 Examples over ℚ(\sqrt{5}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 B.2 Examples over ℚ(\sqrt{2}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 249 B.3 Examples over ℚ(\sqrt{13}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 B.4 Examples over ℚ(\sqrt{3}) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 273

Bibliography 277 iv

Preface

(The “Deutsche Zusammenfassung” following contains the German version of this preface.)

J. Cremona has implemented an algorithm for determining the rank of elliptic curves over ℚ, which is based on 2-descent and described in [B&SD]. Our aim was to develop and implement an analogous algorithm over real quadratic number fields. In 1.3.3 we explain why we have restricted to fields of class number 1.

Up to now, people have only been able to compute ranks over number fields in specific cases. As far as we know, our program is the first general program which – despite timing problems – at least works in principle for elliptic curves over arbitrary real quadratic number fields of class number 1.

This thesis is divided as follows:

Chapter 1 provides basic elements for later chapters. In 1.1 we prove a special version of Hensel’s lemma, which we will need in 2.2 in order to decide whether a quartic equation (or simply a “quartic”) has a solution over a local field. In 1.2 we are dealing with arithmetic in quadratic number fields. We state what the discriminant, the integral basis, the unit group, the decomposition law of rational primes, and the completion with respect to a finite prime look like; how to determine P-adic values; how to find out whether a local or global element is a square and how to compute global square roots; how to determine generators of prime ideals. As mentioned above, we have restricted to fields of class number 1. Therefore we discuss in 1.3 which real quadratic number fields have odd class number and how class number 1 fields are distributed among all real quadratic number fields.

Determining the rank of elliptic curves with non-trivial 2-torsion is based on the so-called 2-descent via 2-isogeny, an algorithm developed by Tate and explained in Chapter 2. In 2.1 we shortly present its main ideas. In 2.2 we describe how to decide whether the quartics occurring are locally soluble everywhere. 2.3 tells how to search for global points on the everywhere locally soluble quartics. The rank can then be deduced from the number of globally soluble quartics. The only problem is that if the program does not find a global point on an everywhere locally soluble quartic, we do not know whether this quartic does not have such a point, or whether we have not searched long enough. From the global points on the quartics we can compute points on the original elliptic curve. 2.4 contains criteria for these points on the elliptic curve to have infinite order and to be linearly independent. (Unfortunately, when we wrote our programs, we did not have a function computing the Néron-Tate height over quadratic fields, which would have solved both problems.) I.e. we do not only obtain the rank r of the elliptic curve but also r linearly independent points. In 2.5 we present two “theoretical tricks” which considerably improve the cpu time of the program for 2-descent via 2-isogeny.

Chapter 3 is about the Tate-Shafarevich group. In 3.1 we list some known results, for example that its order is a square (provided we know that all p-primary components are finite) and that, for a fixed number field K, there exist elliptic curves with arbitrarily large Tate-Shafarevich group over K. 3.2 contains the definition of the ϕ-Selmer and the ϕ-Tate-Shafarevich group for an arbitrary isogeny ϕ. In 3.3 we are dealing with elliptic curves which have a point of order 2, so that the multiplication by 2 on the curve can be written as the composition of an isogeny ϕ and its dual isogeny ϕ̂. We describe how to determine the ϕ- and ϕ̂-Selmer and -Tate-Shafarevich groups in this case. 3.4 concerns the relationship between the ϕ- and ϕ̂-Tate-Shafarevich groups on the one side and the 2-torsion of the Tate-Shafarevich group on the other side. Moreover, we mention how to determine this 2-torsion for elliptic curves with no points of order 2. 3.5 is quite an interesting section. We have tried to construct Tate-Shafarevich groups with large 2-torsion over ℚ(\sqrt{D}), D ∈ \{2, 3, 5, 6, 7, 11, 13, 14, 17, 19\}. And in fact, we found 17 examples with 1024 2-torsion points, which is actually quite remarkable.

Appendix A contains examples of ranks and the 2-torsion of the Tate-Shafarevich group of elliptic curves with a non-trivial point of order 2 over ℚ(\sqrt{5}) and ℚ(\sqrt{6}). In A.1 we present the first four examples we used to test our program for 2-descent via 2-isogeny. They were found by P. Giguere in his Masters thesis at McGill University in Montreal and have rank 0, 1, 2, and 4 over ℚ(\sqrt{5}) and #� (E/ℚ(\sqrt{5}))[2] = 1. In A.2 we have used curves constructed by H. Graf in his Diploma thesis at the Universität des Saarlandes. These curves have rank 4, 5, 6, or 7 over ℚ(\sqrt{5}) or ℚ(\sqrt{6}), and the 2-torsion of the Tate-Shafarevich group has order 1, 4, or 16 in all cases. We have put Appendix A behind Chapter 3 instead of at the end of the thesis, because the examples in Appendix A are supposed to illustrate the theory of Chapter 2 and Chapter 3. (Appendix B contains examples concerning the theory of Chapter 4 and Chapter 5.)

In Chapter 4 we describe general 2-descent, which was developed by Birch and Swinnerton-Dyer and applies to elliptic curves with non-trivial and trivial 2-torsion. In 4.1 we explain the main idea of this method, namely not considering the elliptic curve E itself but quartics whose invariants I and J differ from the invariants c_{4}(E) and 2c_{6}(E) only by a 4th and a 6th power, respectively. In 4.2 we discuss how to determine “all” quartics with invariants I and J for a fixed pair (I, J). In order to obtain a finite number of quartics, we confine ourselves to reduced quartics. Furthermore, we only want one representative for each equivalence class of quartics. For every single step in the search for the quartics, we have tried to compare what had to be done over ℚ on the one side and over a real quadratic number field on the other side. In particular when determining the leading coefficients a of the quartics (4.2.1–4.2.4 and 4.2.6), some additional theoretical work was necessary over real quadratic number fields. Over ℚ we could simply use the usual fundamental domain for the action of SL(2, ℤ) on the complex upper half plane in order to deduce bounds on a. For K = ℚ(\sqrt{D}), D > 0, we first had to find a suitable fundamental domain for the action of SL(2, 𝔒(K)) on the 2-dimensional complex upper half plane; then we had to derive bounds on the norm of a; and finally we had to determine all a whose norm lies between the computed bounds. (Here I would like to thank D. Zagier for his crucial hint.) All other steps (determining all possible types of quartics in 4.2.5, a search region for the coefficient b in 4.2.7, a search region for the coefficient c in 4.2.8, determining the coefficients d and e in 4.2.9) proved to be simpler. But everywhere the “one-dimensional problems” over ℚ became “twodimensional” problems over ℚ(\sqrt{D}), D > 0, because we always had to consider the two possible embeddings of ℚ(\sqrt{D}) into the real numbers. And in general, most of the “little things”, which are simple over ℚ, become more complicated over quadratic number fields. The search region for the second coefficient b of the quartics illustrates this quite well: b must run through a complete set of representatives modulo 4a. Over ℤ we can immediately write down such a set of representatives, namely {b ∈ ℤ; −2|a| < b ≤ 2|a|}. Over the ring of integers of a real quadratic number field, we have to “work” in order to prove that the set of representatives is of the form stated in 4.2.7. For each quartic found we must test whether it belongs to the trivial class (4.3) or whether it is equivalent to a quartic previously obtained, since – as mentioned above – we only need one representative for each equivalence class. Having determined these representatives, we proceed similarly to 2-descent via 2-isogeny: We discard all quartics which are not locally soluble everywhere, search for global points on the remaining quartics (4.5, referring to 2.2 and 2.3), and derive from the number of globally soluble quartics the rank of the elliptic curve. As before we can determine from the global points on the quartics global points on the original elliptic curve (4.6), i.e. we obtain the rank r and r linearly independent points. In 4.7 we show how to speed up the algorithm by using simple congruence conditions (4.7.1) and tricky sieving (4.7.2). Both methods are essential, as – even for curves with small coefficients – the search region for the quartics is so huge that the program would not come to an end in reasonable cpu time if it had to go through the full region.

2-descent (with and without 2-isogeny) is a so-called unconditional algorithm, since it does not depend on any conjecture. But in general the rank can only be determined exactly if we find global points on all everywhere locally soluble quartics, because we usually cannot decide whether an everywhere locally soluble quartic really does not have a global solution – which means that the Tate-Shafarevich group has non-trivial 2-torsion – or whether we have not run the search program long enough.

In the description of Chapter 4, we mentioned that the invariants I and J of the quartics to be determined must differ from c_{4}(E) and 2c_{6}(E) by a 4th and a 6th power, respectively. The question arising in Chapter 5 is which pairs of invariants (I, J) have to be considered. On the one hand, it should be as few pairs as possible, and on the other hand, I and J should be as small as possible so that the search region for the corresponding quartics does not become too large. Hence the main idea is to divide out as many 4th and 6th prime powers from c_{4}(E) and 2c_{6}(E) as possible. In 5.1 we describe how to proceed over ℚ: which conditions imply that c_{4}(E) and 2c_{6}(E) may be replaced by p^{−4} ·c_{4}(E) and p^{−6} ·2c_{6}(E) (see Lemma 3 in 5.1.1 for p ≠ 2, 3, Lemma 4 in 5.1.2 for p = 3, and Lemma 5 in 5.1.3 for p = 2); whether we eventually have to consider one or two pairs of invariants (5.1.4); and how to reduce the cpu time needed for the larger pair considerably in the case of two pairs (5.1.5, 5.1.6). All these results, except those in 5.1.6, were stated by Birch and Swinnerton-Dyer in [B&SD]. But they only gave a brief sketch of the proof of Lemma 3 and omitted the proofs of Lemma 4 and Lemma 5. Our aim in 5.2 was to obtain analogous results for (quadratic) number fields, in particular to find and prove analogous lemmas. The latter proved to be a somewhat thankless task, because it did not require a very high theoretical level – but an enormous amount of time and patience! Lemma 3 over number fields is stated in 5.2.1, Lemma 4 in 5.2.2, Lemma 5^{∗} (a pre-version of Lemma 5) in 5.2.3, and Lemma 5 over quadratic number fields in 5.2.8. 5.2.9 contains the “essence” of Chapter 5: Over quadratic number fields we have to consider at most four pairs of invariants. In 5.2.4–5.2.6 we show, analogously to the case K = ℚ, how to reduce the cpu time needed for the second, third, or fourth pair of invariants if these pairs are relevant.

In Appendix B we have listed examples of ranks of elliptic curves with trivial 2-torsion. Our program runs in reasonable cpu time over ℚ(\sqrt{5}), ℚ(\sqrt{2}), ℚ(\sqrt{13}), and ℚ(\sqrt{3}). We had to construct the examples, because – even over these fields – we cannot control the cpu times of arbitrarily chosen curves, even if these curves have small coefficients, because the search regions for the quartics may become arbitrarily large. However, the current version of the program represents a considerable progress compared to our first version. At the beginning, we were only able to compute ranks over ℚ(\sqrt{5}). Our first example took 24 hours, could then be sped up to 12 hours, and today takes only a few minutes. Moreover, as already mentioned, we are now also able to compute ranks over ℚ(\sqrt{2}), ℚ(\sqrt{13}), and ℚ(\sqrt{3}). And last but not least, one should not forget that this is the first general program for determining ranks over real quadratic number fields of class number 1. The examples in Appendix B have rank 0, 1, 2, or 3. It might be somewhat surprising that rank 0 is relatively rare whereas rank 2 occurs quite often. The reason for this is that the simple curves we chose are rather likely to have simple (linearly independent) points.

The cpu times in Appendix A are much lower than those in Appendix B. This is because in the case of 2-descent via 2-isogeny, the relevant quartics can be determined easily and fast, whereas in the case of general 2-descent, one has to go through a huge search region. It should be pointed out that, contrary to the program for general 2-descent, our program for 2-descent via 2-isogeny works for arbitrary curves over arbitrary real quadratic number fields of class number 1.

All our programs were written in the programming language C (about 15000 lines of code in all), and we ran the examples on the SGI-Challenge “remus” of the Fachbereich Mathematik of the Universität des Saarlandes. In order to determine ranks and the order of the 2-torsion of the Tate-Shafarevich group over ℚ (see 3.5), we used J. Cremonas program mrank in many cases. This program is available free via anonymous ftp from Exeter University. It is based on exactly the same principles as our program over real quadratic number fields (in particular, it is unconditional), namely 2-descent via 2-isogeny for curves with non-trivial 2-torsion and general 2-descent otherwise. We also used functions contained in the computer algebra system simath, which is being developed since 1985 under Prof. Dr. H.G. Zimmer at the Universität des Saarlandes: on the one hand for the computations over ℚ, namely ecrrank (“elliptic curve over the rationals, rank”) and ecrordtsg (”elliptic curve over the rationals, order of the Tate-Shafarevich group”), which are both based on the Birch and Swinnerton-Dyer conjecture (i.e. they are conditional); on the other hand for basic arithmetic in number fields; and finally the multiple precision floating point arithmetic, which has been developed by T. Papanikolaou at the Fachbereich Informatik of the Universität des Saarlandes and is part of the simath system. The reason why we sometimes preferred the conditional simath functions ecrrank and ecrordtsg to J. Cremona’s unconditional mrank, is that we could simply call the simath functions in our C programs and then easily evaluate the output data in the C programs. (simath is also available free via anonymous ftp.)

I would like to thank Prof. Dr. H.G. Zimmer for his advice and his support while I was working on this subject. My very special thanks are due to Dr. J.E. Cremona from Exeter, U.K. My subject came up during a six-month grant at Exeter University, and afterwards my work was based on countless e-mail messages between Saarbrücken and Exeter. This thesis would not have been possible without J. Cremona’s mathematical and moral support. And finally, thank you to all others who have helped – directly or indirectly – to make this thesis possible.

Deutsche Zusammenfassung

J. Cremona hat das in [B&SD] beschriebene, auf 2-Abstieg basierende Verfahren zur Rangberechnung elliptischer Kurven über ℚ implementiert. Ziel unserer Arbeit war es, diesen Algorithmus auf reellquadratische Zahlkörper zu übertragen und zu implementieren. In 1.3.3 ist beschrieben, warum wir uns dabei auf Körper mit Klassenzahl 1 beschränkt haben.

Bisher konnten Ränge über Zahlkörpern nur für spezielle Fälle berechnet werden. Soweit uns bekannt ist, ist unser Programm das erste allgemeine Programm, das prinzipiell – d.h. wenn man davon absieht, daß die Laufzeit beliebig groß werden kann – in der Lage ist, für elliptische Kurven über einem beliebigen reell-quadratischen Zahlkörper mit Klassenzahl 1 den Rang zu bestimmen.

Die Arbeit ist wie folgt aufgebaut:

Kapitel 1 stellt Grundlagen für spätere Kapitel bereit. In 1.1 beweisen wir eine spezielle Version des Hensel’schen Lemmas, die wir in 2.2 brauchen, um zu entscheiden, ob eine quartische Gleichung (eine sogenannte “Quartik”) eine Lösung über einem lokalen Körper hat. In 1.2 ist angegeben, wie für quadratische Zahlkörper die Diskriminante, die Ganzheitsbasis, die Einheitengruppe, das Zerlegungsgesetz rationaler Primzahlen und die vollständige Hülle bzgl. einer endlichen Primstelle aussieht; wie man P-adische Werte bestimmt; wie lokale und globale Quadrate charakterisiert und globale Quadratwurzeln berechnet werden können; und wie man Erzeuger von Primidealen bestimmt. Wie oben schon erwähnt, haben wir uns auf Körper mit Klassenzahl 1 beschränkt. Deshalb haben wir in 1.3 zusammengestellt, welche reell-quadratischen Zahlkörper überhaupt ungerade Klassenzahl haben und wie die Körper mit Klassenzahl 1 unter allen reell-quadratischen Zahlkörpern verteilt sind.

Für elliptische Kurven mit nicht-trivialer 2-Torsion ist es am einfachsten, den Rang zu bestimmen. Das zugrundeliegende, von Tate entwickelte Verfahren, der sog. 2-Abstieg mit 2-Isogenie, ist in Kapitel 2 erläutert. In 2.1 wird der Algorithmus kurz vorgestellt. In 2.2 ist beschrieben, wie man entscheidet, ob die auftretenden Quartiken überall lokal lösbar sind. In 2.3 steht, wie man auf den überall lokal lösbaren Quartiken globale Punkte sucht. Aus der Anzahl der global lösbaren Quartiken ergibt sich dann der Rang. Wenn allerdings das Programm auf einer überall lokal lösbaren Quartik keinen globalen Punkt findet, dann wissen wir nicht, ob diese Quartik keinen solchen Punkt hat oder ob wir nur nicht lange genug gesucht haben. Aus den globalen Punkten auf den Quartiken lassen sich Punkte auf der ursprünglichen elliptischen Kurve konstruieren. 2.4 enthält Kriterien für unendliche Ordnung und lineare Unabhängigkeit dieser Punkte auf der Kurve. (Die Néron-Tate-Höhe über quadratischen Zahlkörpern, mit deren Hilfe beides leicht zu entscheiden wäre, stand uns leider nicht zur Verfügung.) Der 2-Abstieg liefert also nicht nur den Rang r der elliptischen Kurve, sondern auch r linear unabhängige Punkte. In 2.5 stellen wir zwei “theoretische Tricks” vor, die zu einer deutlichen Laufzeitverbesserung führten.

In Kapitel 3 geht es um die Tate-Shafarevich-Gruppe. Zunächst werden in 3.1 bekannte Resultate aufgelistet, wie z.B. daß ihre Ordnung ein Quadrat ist (vorausgesetzt man weiß, daß alle p-primären Komponenten endlich sind) und daß man für jeden festen Zahlkörper K elliptische Kurven mit beliebig großer Tate-Shafarevich-Gruppe über K konstruieren kann. In 3.2 geben wir für eine beliebige Isogenie ϕ die Definition der ϕ-Selmer- und ϕ-Tate-Shafarevich-Gruppe an. In 3.3 gehen wir von einer elliptischen Kurve mit einem nicht-trivialen 2-Teilungspunkt aus, so daß man die Multiplikation mit 2 auf der Kurve aufsplitten kann in eine Isogenie ϕ und die zugehörige duale Isogenie ϕ̂. Wir beschreiben, wie man in diesem Fall die ϕ- und ϕ̂-Selmer- und -Tate-Shafarevich-Gruppe bestimmt. Wie man dann aus den ϕ- und ϕ̂-Tate-Shafarevich-Gruppen Aussagen über die 2-Torsion der Tate-Shafarevich-Gruppe selbst erhält, steht in 3.4. Außerdem wird dort kurz angedeutet, wie man diese 2-Torsion für elliptische Kurven mit nur trivialen 2-Teilungspunkten bestimmt. 3.5 ist ein recht interessanter Abschnitt. Wir haben dort versucht, Tate-Shafarevich-Gruppen mit möglichst vielen 2-Teilungspunkten über ℚ(\sqrt{D}), D ∈ \{2, 3, 5, 6, 7, 11, 13, 14, 17, 19\} zu konstruieren. Und tatsächlich haben wir 17 Beispiele mit 1024 2-Teilungspunkten gefunden, was eigentlich ganz beachtlich ist.

Anhang A enthält Beispiele für Ränge und 2-Torsion der Tate-Shafarevich-Gruppe von elliptischen Kurven mit nicht-trivialen 2-Teilungspunkten über ℚ(\sqrt{5}) und ℚ(\sqrt{6}). In A.1 werden die ersten vier Beispiele vorgestellt, die wir benutzt haben, um unser Programm für 2-Abstieg mit 2-Isogenie zu testen. Sie stammen aus der “Masters thesis” von P. Giguere an der McGill University in Montreal und liefern Rang 0, 1, 2 bzw. 4 über ℚ(\sqrt{5}) und #� (E/ℚ(\sqrt{5}))[2] = 1. In A.2 haben wir Kurven benutzt, die H. Graf in seiner Diplomarbeit an der Universität des Saarlandes konstruiert hat. Diese Kurven haben Rang 4, 5, 6 oder 7 über ℚ(\sqrt{5}) bzw. ℚ(\sqrt{6}), und die 2-Torsion der Tate-Shafarevich-Gruppe hat in allen Fällen Mächtigkeit 1, 4 oder 16. Anhang A steht hinter Kapitel 3 statt am Ende der Arbeit, weil die Beispiele in Anhang A die Theorie aus Kapitel 2 und 3 verdeutlichen sollen. (Anhang B enthält Beispiele zur Theorie in Kapitel 4 und 5.)

In Kapitel 4 beschreiben wir den von Birch und Swinnerton-Dyer entwickelten allgemeinen 2-Abstieg, der auch für elliptische Kurven mit trivialer 2-Torsion angewendet werden kann. In 4.1 wird zunächst die Hauptidee dieses Verfahrens erläutert, nämlich statt der elliptischen Kurve E Quartiken zu betrachten, deren Invarianten I und J sich von den Invarianten c_{4}(E) und 2c_{6}(E) nur um 4. bzw. 6. Potenzen unterscheiden. In 4.2 schildern wir, wie man für ein vorgegebenes Invariantenpaar (I, J) “alle” Quartiken mit diesen Invarianten bestimmt. Damit ihre Anzahl endlich ist, beschränkt man sich auf reduzierte Quartiken. Außerdem interessiert aus jeder Äquivalenzklasse von Quartiken nur ein Vertreter. Wir haben versucht, für jeden einzelnen Schritt bei der Bestimmung der Quartiken die unterschiedlichen Vorgehensweisen über ℚ und über reell-quadratischen Zahlkörpern gegenüberzustellen. Insbesondere bei der Bestimmung der führenden Koeffizienten a der Quartiken (4.2.1–4.2.4 und 4.2.6) war für reell-quadratische Zahlkörper einiges an theoretischem Aufwand nötig. Über ℚ konnte man einfach den üblichen Fundamentalbereich für die Operation von SL(2, ℤ) auf der komplexen oberen Halbebene benutzen, um Schranken für a herzuleiten. Für K = ℚ(\sqrt{D}), D > 0, mußten wir zunächst einen geeigneten Fundamentalbereich für die Operation von SL(2, 𝔒(K)) auf der 2-dimensionalen komplexen oberen Halbebene finden; dann mußten daraus Schranken für die Norm von a abgeleitet werden; und schließlich mußten die zugehörigen a bestimmt werden. (Hier möchte ich D. Zagier für seinen entscheidenden Hinweis danken.) Alle anderen Schritte (Bestimmung der möglichen Typen von Quartiken in 4.2.5, Suchbereich für den Koeffizienten b in 4.2.7, Suchbereich für den Koeffizienten c in 4.2.8, Bestimmung der Koeffizienten d und e in 4.2.9) erwiesen sich als einfacher. Allerdings wurden überall die “ein-dimensionalen Probleme” über ℚ zu “zwei-dimensionalen Problemen” über ℚ(\sqrt{D}), D > 0, weil nämlich immer die beiden möglichen Einbettungen von ℚ(\sqrt{D}) in die reellen Zahlen betrachtet werden mußten. Und ganz allgemein erfordern viele “kleinen Dinge”, die über ℚ einfach sind, über quadratischen Zahlkörpern einiges Nachdenken. Ein gutes Beispiel hierfür ist der Suchbereich für den zweiten Koeffizienten b der Quartiken (4.2.7). b muß ein vollständiges Repräsentantensystem modulo 4a durchlaufen. Über ℤ läßt sich ein solches Repräsentantensystem sofort hinschreiben, nämlich {b ∈ ℤ; −2|a| < b ≤ 2|a|}. Über reell-quadratischen Zahlkörpern aber muß man “arbeiten”, bis man gezeigt hat, daß das Repräsentantensystem die in 4.2.7 angegebene Form hat. Für jede gefundene Quartik muß man testen, ob sie zur trivialen Äquivalenzklasse gehört (4.3) und ob sie zu einer der vorher gefundenen Quartiken äquivalent ist (4.4), denn wie oben erwähnt brauchen wir aus jeder Äquivalenzklasse nur einen Vertreter. Wenn man diese Vertreter bestimmt hat, geht man ähnlich vor wie beim 2-Abstieg mit 2-Isogenie: Man wirft alle Quartiken weg, die nicht überall lokal lösbar sind, sucht auf den verbleibenden Quartiken globale Punkte (4.5, bzw. 2.2 und 2.3), und erhält aus der Anzahl der global lösbaren Quartiken den Rang der elliptischen Kurve. Auch hier lassen sich aus den globalen Punkten auf den Quartiken Punkte auf der ursprünglichen elliptischen Kurve bestimmen (4.6), d.h. man erhält auch hier den Rang r und r linear unabhängige Punkte. In 4.7 zeigen wir, wie man mit einfachen Kongruenzbetrachtungen (4.7.1) und geschicktem Siebverfahren (4.7.2) den Algorithmus beschleunigen kann. Beide Methoden sind eigentlich unerläßlich, denn der Suchbereich für die Quartiken ist selbst für Kurven mit kleinen Koeffizienten so riesig, daß das Programm nicht in vernünftiger Zeit enden würde, wenn man diesen Suchbereich komplett durchlaufen würde.

2-Abstieg (mit und ohne 2-Isogenie) ist ein sog. “unconditional algorithm”, ein Verfahren, das von keiner Vermutung abhängt. Allerdings kann man im allgemeinen nur dann den Rang exakt bestimmen, wenn man auf allen überall lokal lösbaren Quartiken globale Punkte findet. Denn man kann wie gesagt ohne theoretischen Zusatzaufwand i.a. nicht entscheiden, ob eine überall lokal lösbare Quartik wirklich keine globale Lösung besitzt – was bedeutet, daß die Tate-Shafarevich-Gruppe nicht-triviale 2-Torsion hat – oder ob das Suchprogramm nicht lange genug gelaufen ist.

In der Beschreibung von Kapitel 4 hatten wir erwähnt, daß sich die Invarianten I und J der zu bestimmenden Quartiken von c_{4}(E) und 2c_{6}(E) um eine 4. bzw. eine 6. Potenz unterscheiden müssen. In Kapitel 5 stellt sich nun die Frage, welche Invariantenpaare (I, J) man betrachten muß. Zum einen sollten es möglichst wenige Paare sein, und zum andern sollten I und J möglichst klein sein, um den Suchbereich für die zugehörigen Quartiken in vernünftigen Grenzen zu halten. Also ist die Grundidee zu versuchen, aus c_{4}(E) und 2c_{6}(E) möglichst viele 4. bzw. 6. Primpotenzen herauszudividieren. In 5.1 beschreiben wir, wie man über ℚ vorgeht: unter welchen Bedingungen man c_{4}(E) bzw. 2c_{6}(E) durch p^{−4}·c_{4}(E) bzw. p^{−6}·2c_{6}(E) ersetzen darf (für p ≠ 2, 3 in Lemma 3 in 5.1.1, für p = 3 in Lemma 4 in 5.1.2 und für p = 2 in Lemma 5 in 5.1.3); ob man letztendlich ein oder zwei Invariantenpaare betrachten muß (5.1.4); und wie man, falls man zwei Paare betrachten muß, die Laufzeit für das größere Paar deutlich reduzieren kann (5.1.5, 5.1.6). Alle diese Ergebnisse (außer 5.1.6) stammen von Birch und Swinnerton-Dyer, die aber in [B&SD] nur ganz kurz den Beweis von Lemma 3 skizziert haben und die Beweise von Lemma 4 und Lemma 5 weggelassen haben. Unser Ziel in 5.2 war es, analoge Resultate für (quadratische) Zahlkörper zu erhalten, insbesondere analoge Lemmata zu finden und zu beweisen. Letzteres erwies sich als etwas undankbare Aufgabe, weil das theoretische Niveau nicht besonders hoch war, der Aufwand an Zeit und Geduld aber enorm! Lemma 3 über Zahlkörpern steht in 5.2.1, Lemma 4 in 5.2.2, Lemma 5^{∗} (eine Vorversion zu Lemma 5) in 5.2.3 und Lemma 5 über quadratischen Zahlkörpern in 5.2.8. 5.2.9 enthält die Quintessenz von Kapitel 5, daß man nämlich über quadratischen Zahlkörpern höchstens vier Invariantenpaare betrachten muß. In 5.2.4–5.2.6 ist beschrieben, wie man, analog zur Rangberechnung über ℚ, ggfs. für das zweite, dritte und vierte Invariantenpaar die Laufzeit verringern kann.

In Anhang B haben wir Beispiele für Ränge elliptischer Kurven mit trivialer 2-Torsion angegeben.

Unser Programm läuft in vernünftiger Zeit für Kurven über ℚ(\sqrt{5}), ℚ(\sqrt{2}), ℚ(\sqrt{13}) und ℚ(\sqrt{3}). Die angebenen Beispiele mußten wir konstruieren, weil auch über diesen Körpern die Laufzeiten nicht in vertretbarem Rahmen zu halten sind, wenn man beliebige Kurven nimmt, selbst wenn diese kleine Koeffizienten haben, denn der Suchbereich für die Quartiken kann beliebig groß werden. Trotzdem stellt die jetzige Version des Programms einen erheblichen Fortschritt gegenüber der ersten lauffähigen Version dar: Anfangs konnten wir nur Ränge über ℚ(\sqrt{5}) berechnen. Unser erstes Beispiel lief 24 Stunden, konnte dann auf 12 Stunden verbessert werden und braucht heute nur noch wenige Minuten. Mittlerweile sind, wie schon erwähnt, auch Rangberechnungen über ℚ(\sqrt{2}), ℚ(\sqrt{13}) und ℚ(\sqrt{3}) möglich. Außerdem sollte man nicht vergessen, daß unser Programm das erste allgemeine Programm ist, das Ränge über reell-quadratischen Zahlkörpern mit Klassenzahl 1 bestimmen kann. Die Beispiele in Anhang B haben Rang 0, 1, 2 oder 3. Auffallend ist, daß Rang 0 relativ selten und Rang 2 relativ oft vorkommt. Der Grund dafür ist, daß für diese einfachen Kurven die Wahrscheinlichkeit, daß sie einfache (linear unabhängige) Punkte haben, ziemlich hoch ist.

Die Laufzeiten in Anhang A sind deutlich niedriger als die in Anhang B. Das liegt daran, daß beim 2-Abstieg mit 2-Isogenie die relevanten Quartiken ganz einfach bestimmt werden können, während beim allgemeinen 2-Abstieg ein riesiger Suchbereich durchlaufen werden muß. Im Gegensatz zu dem Programm für allgemeinen 2-Abstieg läuft das Programm für 2-Abstieg mit 2-Isogenie für beliebige Kurven über beliebigen reell-quadratischen Zahlkörpern mit Klassenzahl 1.

Alle unsere Programme sind in der Programmiersprache C geschrieben (insgesamt etwa 15000 Zeilen Programm-Code), und die Beispiele liefen auf der SGI-Challenge “remus” des Fachbereichs Mathematik der Universität des Saarlandes. Die notwendigen Rechnungen über ℚ (d.h. Bestimmung von Rang und 2-Torsion der Tate-Shafarevich-Gruppe, siehe 3.5) wurden zum großen Teil mit J. Cremonas Programm mrank durchgeführt. Dieses Programm ist über “anonymous ftp” von der Universität Exeter erhältlich und beruht auf denselben Prinzipien wie unser Programm über reell-quadratischen Zahlkörpern (ist also insbesondere “unconditional”), nämlich 2-Abstieg mit 2-Isogenie für Kurven mit nicht-trivialen 2-Teilungspunkten und allgemeiner 2-Abstieg sonst. Außerdem haben wir Funktionen des Computeralgebra-Systems simath benutzt, das seit 1985 am Lehrstuhl von Prof. Dr. H.G. Zimmer an der Universität des Saarlandes entwickelt wird; zum einen für die Rechnungen über ℚ, nämlich ecrrank (“elliptic curve over the rationals, rank”) und ecrordtsg (“elliptic curve over the rationals, order of the Tate-Shafarevich group”), die beide auf der Vermutung von Birch und Swinnerton-Dyer basieren (also “conditional” sind); zum anderen für die Grundarithmetik in Zahlkörpern; und schließlich noch die von T. Papanikolaou am Fachbereich Informatik der Universität des Saarlandes entwickelte beliebig genaue reelle Arithmetik, die in das simath-System eingebaut ist. Der Grund, warum wir manchmal die “conditional” simath-Funktionen ecrrank und ecrordtsg J. Cremonas “unconditional” mrank vorgezogen haben, ist, daß man simath-Funktionen einfach in seinen C-Programmen aufrufen kann, und sich dann auch die Ausgabedaten leicht in den C-Programmen auswerten lassen. (simath ist übrigens auch über “anonymous ftp” erhältlich.)

An dieser Stelle möchte ich mich bei Prof. Dr. H.G. Zimmer bedanken für seine Anregungen und seine Unterstützung bei der Ausarbeitung dieses Themas. Mein ganz besonderer Dank gilt Dr. J.E. Cremona aus Exeter, Großbritannien. Mein Thema ist während eines sechsmonatigen Stipendiums an der Universität Exeter entstanden, und die Ausarbeitung wurde danach durch unzählige e-mails zwischen Saarbrücken und Exeter vorangetrieben. Ohne J. Cremonas fachliche und auch moralische Unterstützung wäre diese Arbeit nicht möglich gewesen. Und nicht zuletzt möchte ich auch allen anderen, die direkt oder indirekt zum Gelingen der Arbeit beigetragen haben, herzlich danken.

Bibliography

[B&SD] B.J. Birch and H.P.F. Swinnerton-Dyer, Notes on elliptic curves. I., J. Reine Angew. Math.

212 (1963), 7–25

[Ca] J.W.S. Cassels, Arithmetic on curves of genus 1 (IV). Proof of the Hauptvermutung, J. Reine Angew. Math. 211 (1962), 95–112

[Co-Le] H. Cohen and H.W. Lenstra, Heuristics on class groups of number fields, Lecture Notes in

Math. 1068 (1983), 33–62

[Co] H. Cohn, A second course in number theory, Wiley, 1962

[Con] I. Connell, Addendum to a paper of Harada and Lang, J. Algebra 145 (1992), 463–467 [Cox] D.A. Cox, Primes of the form x^{2} + ny^{2}, Wiley, 1989

[Cr] J.E. Cremona, Algorithms for modular elliptic curves, Cambridge University Press, 1992 [Fr] G. Frey, Some aspects of the theory of elliptic curves over number fields, Exposition. Math. 4

(1986), 35–66

[Ga] P.B. Garrett, Holomorphic Hilbert modular forms, Wadsworth, 1990

[Graf] H. Graf, Konstruktion elliptischer Kurven hohen Ranges über quadratischen Zahlkörpern der Klassenzahl eins, Diploma thesis, Universität des Saarlandes, Saarbrücken, 1995

[Gr] C. Grasser, Lokale Invarianten einer elliptischen Kurve E/K, Diploma thesis, Universität des Saarlandes, Saarbrücken, 1992

[Ha] H. Hasse, Number theory, Springer-Verlag, 1980

[Hu] D. Husemöller, Elliptic curves, Springer-Verlag, 1987 [Ir-Ro] K. Ireland and M. Rosen, A classical introduction to modern number theory, Springer-Verlag,

1980

[Ka] S. Kamienny, Torsion points on elliptic curves, Bull. Amer. Math. Soc. 23 (1990), 371–373 [Ke-Mo] M.A. Kenku and F. Momose, Torsion points on elliptic curves defined over quadratic fields,

Nagoya Math. J. 109 (1988), 125–149

[Ko] N. Koblitz, Introduction to elliptic curves and modular forms, Springer-Verlag, 1984 [Ko-Pi] H. Koch und H. Pieper, Zahlentheorie, Deutscher Verlag der Wissenschaften, 1976

[Kol] V.A. Kolyvagin, Finiteness of E(ℚ) and � (E/ℚ) for a subclass of Weil curves, Math. USSR Izv. 32 (1989), 523-542 278 BIBLIOGRAPHY

[Kr] K. Kramer, A family of semistable elliptic curves with large Tate-Shafarevitch groups, Proc. Amer. Math. Soc. 89, 3 (1983), 379–386

[Ma] J. Martinet, Propriétés arithmétiques des cubiques de genre un dans un corps de nombres algébriques fini, PhD thesis, Université de Grenoble, 1964

[Mo-Wi] R.A. Mollin and H.C. Williams, Computation of the class number of a real quadratic field,

preprint

[Pe] A. Pethö, Algebraische Algorithmen, Vieweg-Verlag, 1996, to appear

[Ru] K. Rubin, Tate-Shafarevich groups and L-functions of elliptic curves with complex multiplication, Invent. Math. 89 (1987), 527–559

[Sik] S. Siksek, Descents on curves of genus 1, PhD thesis, Exeter University, 1995

[Si] J.H. Silverman, The arithmetic of elliptic curves, Springer-Verlag, 1986 [Si-Ta] J.H. Silverman and J. Tate, Rational points on elliptic curves, Springer-Verlag, 1992

[Ta] J. Tate, Rational points on elliptic curves, Lectures held at Haverford College, 1961

[We] E. Weiss, Algebraic number theory, McGraw-Hill, 1963

[Wi] H.C. Williams, A numerical investigation into the length of the period of the continued fraction expansion of \sqrt{D}, Math. Comp. 36, 154 (1981), 593–601

[Zi] H.G. Zimmer, Torsion points on elliptic curves over a global field, Manuscripta Math. 29 (1979), 119–145